# Flavored QCD axion and Modular invariance

### v2

@@ Appendix A The potentials $V^{(0)}_{F}$ and $V^{(1)}_{F}$
$V^{(1)}_{F}=e^{K/M^{2}_{P}}\frac{C_{0}M^{6}_{P}}{|\eta(\tau)|^{4h}}\Big\{\big(-D(\bar{S}+\frac{h}{2\pi^{2}}\ln\eta(\bar{\tau}))-{\rm h.c.}\big)\Big(K^{\tau\bar{\tau}}|H|^{2}+\frac{K^{S\bar{S}}}{y^{2}}\Big)$
$\qquad\qquad-\frac{2\sigma}{M^{2}_{P}}(Aae^{-a\sigma}-Bbe^{-b\sigma})(S+\bar{S}+\frac{h}{2\pi^{2}}\ln\eta(\bar{\tau})\eta(\tau))+\frac{K^{S\bar{S}}}{y}(D+\bar{D})$
$\qquad\qquad+\frac{h}{2\pi^{2}}K^{\tau\bar{\tau}}\Big(DH\frac{\eta^{\prime}(\bar{\tau})}{\eta(\bar{\tau})}+\bar{D}\bar{H}\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\Big)\Big\}\,.$ (108)
@@ Appendix B SUSY breaking in the direction of $U_{X}$
$m^{2}_{\sigma}=\frac{\sigma_{0}}{9\langle y\rangle(2{\rm Im}\tau_{0})^{3}}\Big|A_{0}a^{2}\Big(\frac{aA_{0}}{bB_{0}}\Big)^{\frac{a}{b-a}}+B_{0}b^{2}\Big(\frac{aA_{0}}{bB_{0}}\Big)^{\frac{b}{b-a}}\Big|^{2}\frac{M^{2}_{P}}{|\eta(\tau_{0})|^{12}}\,.$ (109)
@@
where $\Delta W=\langle W\rangle_{\rm AdS}$ is the value of the superpotential at the AdS minimum. By introducing an uplifting potential of the form $\Delta V\approx|V_{\rm AdS}|(\sigma_{0}/\sigma)^{n}$, where typical choices are $n=2$ (as in KKLT Kachru:2003sx ) or $n=3$ (as in Ref. Burgess:2003ic ), the AdS minimum turns into a dS minimum. This uplifting simultaneously breaks supersymmetry. The uplifting induces a shift $\Delta\sigma=\sigma-\sigma_{0}$ in the $U_{X}$ field, see Eq.(32), determined by solving the minimization condition $\partial_{\sigma}(V+\Delta V)|_{\sigma_{0}+\delta\sigma}=0$ along the $\sigma={\rm Re}[U_{X}]$ direction. For the case of a D-term induced uplift parametrized as $\Delta V=\frac{1}{2}(\xi^{\rm FI}_{X})^{2}g^{2}_{X}\approx|V_{\rm AdS}|(\sigma_{0}/\sigma)^{3}$Burgess:2003ic , and using the approximation $(D_{\sigma}W)_{\sigma}\approx W_{\sigma\sigma}|_{\sigma_{0}}$, we obtain the displacement $\Delta\sigma\simeq\langle y\rangle(2{\rm Im}\,\tau_{0})^{3}18M^{2}_{P}|V_{\rm AdS}|/(W_{\sigma\sigma})^{2}$. This leads to the SUSY-breaking $F$-term in the uplifted minimum $D_{\sigma}W\simeq(D_{\sigma}W)_{\sigma}\Delta\sigma\simeq W_{\sigma\sigma}(\sigma_{0})\Delta\sigma$:
@@ Appendix C Scalar potential for flavon
In the global SUSY limit, i.e. $M_{P}\rightarrow\infty$, the scalar potential is determined by the vanishing of the $F$- and $D$-terms for all fields. Then the relevant $F$-term from Eq.(47) and $D$-term in Eq.(29) scalar potentials reads
$V_{F}^{\rm global}$ $=$ $\big|g_{\chi}\chi\tilde{\chi}-\mu^{2}_{\chi}\big|^{2}\,,\qquad V_{D}^{\rm global}=\frac{g^{2}_{X}}{2}\big(-\xi^{\rm FI}_{X}+|\chi|^{2}-|\tilde{\chi}|^{2}\big)^{2}\,\,.$ (112)
The scalar fields $\chi$ and $\tilde{\chi}$, with $X$-charges +1+1 and 1-1 respectively, transform under $U(1)_{X}$ as
$\chi\rightarrow e^{+i\xi}\chi\,,\qquad\tilde{\chi}\rightarrow e^{-i\xi}\tilde{\chi}\,$, (113)
with a constant $\xi$, ensuring that the potential $V_{\rm SUSY}$ respects $U(1)_{X}$ symmetry. Since supersymmetry is preserved after spontaneous breaking of $U(1)_{X}$, the scalar potential vanishes at its ground states, i.e. $\langle V_{F}^{\rm global}\rangle=0$ and $\langle V_{D}^{\rm global}\rangle=0$. A vanishing $F$-term implies a vanishing $D$-term as well.

### v3

@@ Appendix A The potentials $V^{(0)}_{F}$ and $V^{(1)}_{F}$
$V^{(1)}_{F}=e^{K/M^{2}_{P}}\frac{C_{0}M^{6}_{P}}{|\eta(\tau)|^{4h}}\Big\{\big(-D(\bar{S}+\frac{h}{8\pi^{2}}\ln\eta(\bar{\tau}))-{\rm h.c.}\big)\Big(K^{\tau\bar{\tau}}|H|^{2}+\frac{K^{S\bar{S}}}{y^{2}}\Big)$
$\qquad\qquad-\frac{2\sigma}{M^{2}_{P}}(Aae^{-a\sigma}-Bbe^{-b\sigma})(S+\bar{S}+\frac{h}{8\pi^{2}}\ln\eta(\bar{\tau})\eta(\tau))+\frac{K^{S\bar{S}}}{y}(D+\bar{D})$
$\qquad\qquad+\frac{h}{8\pi^{2}}K^{\tau\bar{\tau}}\Big(DH\frac{\eta^{\prime}(\bar{\tau})}{\eta(\bar{\tau})}+\bar{D}\bar{H}\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\Big)\Big\}\,.$ (109)
@@ Appendix B SUSY breaking in the direction of $U_{X}$
$m^{2}_{\sigma}=\frac{\sigma_{0}}{9\langle y\rangle(2{\rm Im}\tau_{0})^{3}}\Big|A_{0}a^{2}\Big(\frac{aA_{0}}{bB_{0}}\Big)^{\frac{a}{b-a}}-B_{0}b^{2}\Big(\frac{aA_{0}}{bB_{0}}\Big)^{\frac{b}{b-a}}\Big|^{2}\frac{M^{2}_{P}}{|\eta(\tau_{0})|^{12}}\,.$ (110)
@@
where $\Delta W=\langle W\rangle_{\rm AdS}$ is the value of the superpotential at the AdS minimum. By introducing an uplifting potential of the form $\Delta V\approx|V_{\rm AdS}|(\sigma_{0}/\sigma)^{n}$, where typical choices are $n=2$ (as in KKLT Kachru:2003sx ) or $n=3$ (as in Ref. Burgess:2003ic ), the AdS minimum turns into a dS minimum. This uplifting simultaneously breaks supersymmetry. The uplifting induces a shift $\Delta\sigma=\sigma-\sigma_{0}$ in the $U_{X}$ field, see Eq.(33), determined by solving the minimization condition $\partial_{\sigma}(V+\Delta V)|_{\sigma_{0}+\delta\sigma}=0$ along the $\sigma={\rm Re}[U_{X}]$ direction. For the case of a D-term induced uplift parametrized as $\Delta V=\frac{1}{2}(\xi^{\rm FI}_{X})^{2}\tilde{g}^{2}_{X}\approx|V_{\rm AdS}|(\sigma_{0}/\sigma)^{3}$Burgess:2003ic , and using the approximation $(D_{\sigma}W)_{\sigma}\approx W_{\sigma\sigma}|_{\sigma_{0}}$, we obtain the displacement $\Delta\sigma\simeq\langle y\rangle(2{\rm Im}\,\tau_{0})^{3}18M^{2}_{P}|V_{\rm AdS}|/(W_{\sigma\sigma})^{2}$. This leads to the SUSY-breaking $F$-term in the uplifted minimum $D_{\sigma}W\simeq(D_{\sigma}W)_{\sigma}\Delta\sigma\simeq W_{\sigma\sigma}(\sigma_{0})\Delta\sigma$:
@@ Appendix C Scalar potential for flavon
Setting to zero from the beginning the matter fields $\{q^{c},\ell,H_{u},$...\}, with the almost vanishing cosmological constant for the remaining fields the gravitino mass $m_{3/2}$ is directly related to the scale of supersymmetry breaking $V=|F|^{2}-3m^{2}_{3/2}M^{2}_{P}+\frac{1}{2}|D|^{2}\approx 0$ implying that the $F$, $D$-terms should vanish in the limit $m_{3/2}$ going to zero and the $F$-term should scale as $m_{3/2}$ at the minimum for the vanishing $D$-term.
As shown in Sec. II.3, the leading-order VEV of $U_{X}$ is fixed at a supersymmetric Minkowski minimum, see Eq.(38), where the $S$-dependence part of $W(S,U_{X},\tau)$ decouples in the limit $\alpha\rightarrow 0$. At this point the modulus $U_{X}$ is strongly stabilized supersymmetrically at $\sigma_{0}$Kachru:2003aw . In this supersymmetric limit, the vacuum configuration of the fields $\chi,\tilde{\chi},\chi_{0}$ is determined by requiring the $F$-terms to vanish, that is, $D_{\chi}W=D_{\tilde{\chi}}W=D_{\chi_{0}}W=0$, where $W=W_{v}+W(U_{X},\tau)$ from Eqs.(34) and (48). Explicitly,
$D_{\chi}W=g_{\chi}\chi_{0}\tilde{\chi}+\frac{W}{M^{2}_{P}}\tilde{\chi}\,$,
$D_{\tilde{\chi}}W=g_{\chi}\chi_{0}\chi+\frac{W}{M^{2}_{P}}\chi\,$,
$D_{\chi_{0}}W=g_{\chi}\chi\tilde{\chi}-\mu^{2}_{\chi}+\frac{W}{M^{2}_{P}}(-i\tau+i\bar{\tau})^{-h}\bar{\chi}_{0}\,.$ (113)
From this set of three equations, we can obtain the supersymmetric vacuum for $\chi,\tilde{\chi},\chi_{0}$
$\langle\chi_{0}\rangle=0\,,\qquad\langle\chi\rangle=\langle\tilde{\chi}\rangle=\frac{v_{\chi}}{\sqrt{2}}\quad\text{with}~\mu_{\chi}=v_{\chi}\sqrt{\frac{g_{\chi}}{2}}\,$, (114)
which satisfies $\langle W\rangle=0$ in the supersymmetric limit due to Eq.(38).
Next, the $D$-term potential is included as a small perturbation $\Delta V$, see above Eq.(112). The $D$-term potential, see Eq.(29) and above Eq.(33), reads
$V_{D}=\frac{\tilde{g}^{2}_{X}}{2}\big(-\xi^{\rm FI}_{X}+|\chi|^{2}-|\tilde{\chi}|^{2}\big)^{2}\,$, (115)
where $|\chi|^{2}-|\tilde{\chi}|^{2}$ corresponds to $X|\varphi_{X}|^{2}$ in the expression for $V_{D}$ above Eq.(33). The scalar fields $\chi$ and $\tilde{\chi}$, carrying $U(1)_{X}$ charges +1+1 and 1-1, respectively, transform as $\chi\rightarrow e^{+i\xi}\chi$ and $\tilde{\chi}\rightarrow e^{-i\xi}\tilde{\chi}$ with a constant $\xi$, ensuring that the potential $V_{\rm SUSY}$ respects $U(1)_{X}$ symmetry. The $D$-term constraint $|\chi|^{2}-|\tilde{\chi}|^{2}-\xi^{\rm FI}_{X}(U_{X})=0$ becomes a small shift condition on $\sigma$ around $\sigma_{0}$, inducing only small subleading corrections to the $F$-term minimum Eq.(B3). Since the $F$-term potential dominates the stabilization (the mass of $\sigma$ from $V_{F}$ is large Eq.(B1)), then the $D$-term just shifts $\sigma_{0}$ slightly, see above Eq.(B2), and the vacuum becomes non-supersymmetric with small positive energy (uplift to dS space).
Finally, since all the soft SUSY breaking parameters in $V_{\rm soft}$ are expected to be of order $m_{3/2}$ which is much smaller than the mass scales appearing in $W_{v}$, it makes sense to minimize the scalar potential $V_{\rm total}$ in the supersymmetric limit and subsequently incorporate the soft breaking effects as small corrections.